OpenStudy (gigi1103):

If you can read 40 pages in 2 hours, how many pages can you read in 5 hours?

OpenStudy (gigi1103):

@AdamK

OpenStudy (anonymous):

100

OpenStudy (anonymous):

100 pages

OpenStudy (gigi1103):

helllllpppppppppppp i give medals

OpenStudy (anonymous):

350 i think

OpenStudy (anonymous):

100 nvm

OpenStudy (adamk):

Set up a proportion. 40/2 and x/5 Cross multiply to get: 2x = 200 x = 100

OpenStudy (gigi1103):

can you guys explain

OpenStudy (gigi1103):

oh thanks adamk

OpenStudy (anonymous):

it is 100

OpenStudy (kendricklamar2014):

40 pages in 2 hours 80 pages in 4 hours 20 pages in 1 hour 100 pages in 5 hours @gigi1103

OpenStudy (adamk):

@C3L3STA Please don't just give answers.

OpenStudy (anonymous):

yea ... ok you can read 40 pages in 2 hour. divide the number of pages by the number of hours. you get 40 / 2 = 20 so now you know you can read 20 pages per hour. they asked you how many in 5 hour. no problem. just do times 5. 5 * 20 = 100

OpenStudy (gigi1103):

omg i get it

OpenStudy (gigi1103):

let me fan you guys

OpenStudy (gigi1103):

sorry my computers slow but hold on

OpenStudy (gigi1103):

one more plz

OpenStudy (gigi1103):

Mark $100 for 10 hours Rick $80 for 9 hours Steve $75 for 8 hours Andy $60 for 5 hours Meg wants to take tennis lessons. In the chart, four instructors are listed with what they charge. Which instructor gives the best deal? A) Mark B) Rick C) Andy D) Steve

OpenStudy (gigi1103):

brb

OpenStudy (gigi1103):

@jpgonzales55

OpenStudy (gigi1103):

@ElonaSushchik

OpenStudy (adamk):

Please make a new question.

OpenStudy (gigi1103):

ok

OpenStudy (anonymous):

C

OpenStudy (anonymous):

its c